# AttributionScanner: A Visual Analytics System for Model Validation with Metadata-Free Slice Finding

### v2

@@ III-F System Design Considerations
III-G Slice Error Mitigation
@@ IV-A Hair Color Classifier Validation - Finding Edge Cases
$\bullet$ Does the model behave correctly on well-performing slices? One basic expected behavior for a hair color classification model is to catch hair features. At first glance of the Attribution Mosaic (Fig. 2  C), ML experts notice slice_22 and slice_5 on the left, which lie separately with others on the Attribution Mosaic. Their mosaic visualizations indicate gray hair patterns, which means the model uses the correct features, i.e., core features. By verifying the corresponding attribution heatmaps (Fig. 2  D), they confirm the correctness of this insight and annotate both slices as core feature with description Core: gray hair. Upon experts saving the annotation, AttributionScanner automatically propagates the annotation and provides a Spuriousness probability of each slice. With this guidance, the experts observe many slices lying on the right of Attribution Mosaic have higher Spuriousness probabilities (Fig. 2  C  E). Their mosaic visualizations do not show any hair patterns, leading to valid doubts that the model does not behave correctly on such slices. Besides, they also notice that the model has correct predictions on these slices (with $100\%$ prediction accuracy). This makes them worry that the model is largely biased by spurious features. Through investigation, they find the model mistakenly utilizes mouth and eyes to predict hair color for those top-performed slices, such as slice_25, slice_14, and slice_2 (Fig. 9).
$\bullet$ Why slices underperform? The experts are also eager to investigate the underperformed slices. By sorting the Slice Table (Fig. 2  B) by ascending order of accuracy, they notice the model only achieves $72.41\%$ accuracy on slice_44 and find that the slice’s visualization only shows colorful patterns without any recognizable content. After checking the attribution heatmaps, they find that the model mistakenly uses image backgrounds to make hair color predictions (Fig. 9). This spurious correlation issue stands out, and they annotate this slice as “spurious” with the description “Spurious: backgrounds”. Similar issues also occur in its neighborhood slices via Attribution Mosaic, and AttributionScanner automatically propagates them with higher Spuriousness.
$\bullet$ What underlying factors contribute to unexpected behaviors? ML experts are interested in understanding why such unexpected model behaviors happen. By sorting the Slice Table by descending order of Spuriousness, they get a list of slices with high Spuriousness possibilities. They switch Attribution Mosaic into the confusion matrix form (Fig. 2  E) to study further details. By clicking on the name of “slice_35”, they highlight this slice on the four sub-views and inspect the provided explanations, where they notice that images that lie in the “FN” group of this slice have wrong labels — they should be labeled as “not gray hair” rather than “gray hair”. They mark this issue as wrong labels and wonder whether this situation is common. By investigating its neighborhood slices on Attribution Mosaic, they find and mark another slice with wrong labels through a few clicks, slice_2_FN.
@@ IV-B Bird Category Classifier Validation - Detecting Bias
$\bullet$ Does the model exhibit bias? To answer this key question, experts start by investigating the underperformed slices. From the Slice Table (Fig. 6  A), they sort slices by ascending order of accuracy and select the worst-performed slice_38. The coordinated information provided by Attribution Mosaic and model attribution heatmaps highlights a spurious correlation problem—the model uses water backgrounds to classify birds (Fig. 10). The experts annotate this slice as “spurious”, and AttributionScanner automatically propagates this annotation. They verify the propagation correctness on neighborhood slices and annotate slice_41 as “spurious” with “water backgrounds” (Fig. 10). Moreover, the experts identify an underperformed slice_32 that is not clustered with the current ones and is given a high Spuriousness possibility. They investigate it and verify that the model utilizes “Spurious feature: land backgrounds”, to predict bird classes. Through the Attribution Mosaic in the confusion matrix form (Fig. 6  D), they find such spurious correlations result in many false negatives (Fig. 6  D1), where the model uses land backgrounds to mistakenly predict many “waterbirds” as “landbirds”.
@@
$\bullet$ Is the detected bias prevalent across all slices? Why or why not? ML experts are interested in determining whether the detected bias is pervasive throughout the dataset. By analyzing slices that are distant from the annotated ones in the Attribution Mosaic and are assigned with low Spuriousness possibilities, they discovered that the farthest neighbors, namely slice_34 and slice_5, correspond to core features. This suggests that the model can correctly capture bird regions in these slices (Fig. 6  B  C), which raises a follow-up “why” question. To understand in what circumstances when the model fails. They browse the original images from slice_32 (spurious feature) and slice_34 (core feature), respectively, as shown in Fig. 6  D and  C. They find slice_32 has very similar land backgrounds and very different birds, while on the other hand, the birds’ appearance of slice_34 is very consistent. This finding explains why this biased model can successfully capture the core features from slice_34 but fails at slice_32—greater similarities in the representation space indicate stronger features. In other words, core features in slice_42 are strong enough to support the model’s robustness against bias. Such insights are helpful in improving model robustness and have been further studied by ML experts [43].
@@
V Evaluation
V-A Quantitative Evaluation
$\bullet$ RCS. This matrix quantifies the model’s reliance on core features while controlling for general noise robustness. It is defined as the ratio of the absolute gap between core and spurious accuracy, and the total possible gap for any model between core and spurious accuracy. Represented as $RCS=\frac{acc^{(C)}-acc^{(S)}}{2\times min(\alpha,1-\alpha)}$, where $\alpha=\frac{acc^{(C)}+acc^{(S)}}{2}$. RCS ranges from 00 to $1$, a higher value indicative of better model performance.
@@ V-C Experts Feedback
VI Discussion and Future Work
@@
Propagation approaches for the hypothetical spuriousness.     We adopt the Label Propagation method [39] to automatically spread the slices’ spurious annotations from users to other slices as discussed in Sec. III-E. This is under the assumption that the distance between the slices corresponding to spurious correlations (e.g., water backgrounds) and the slices corresponding to core correlations (e.g., birds) is large. However, such assumptions can be invalid in other scenarios. For example, an image classification model trained on ImageNet [52] includes various different classes such as “husky dog” and “king penguin”. It is likely that two slices far away from each other are both corresponding to the core correlations. To address this, we plan to incorporate class similarities to the Label Propagation matrix to improve the precision of the generated hypothetical Spuriousness.
@@
VII Conclusion
In this work, we present AttributionScanner, a novel human-in-the-loop model validation system empowered by metadata-free data slice finding and XAI techniques, which can support effective detection, explanation, and mitigation of potential errors and biases in vision models. To achieve this, we first present a metadata-free data slice finding method by using attribution-weighted feature vectors to group unstructured image data into data slices with consistent model attributions. Also, we design a slice pattern summarization view, Attribution Mosaic, which visually presents the shared significant patterns of each slice attributing to the model’s decision. This way, AttributionScanner is able to provide explainable data slices that reveal critical model problems, such as spurious correlations and mislabeled data. We also demonstrate the effectiveness of this work in performance improvement and bias mitigation through both quantitative and qualitative evaluations. To sum up, without additional meta-information, multi-modal model, or other background information, AttributionScanner enables an effective model validation and error mitigation workflow to alleviate underlying issues in ML models. We hope this work can inspire future research in human-AI teaming to improve AI trustworthiness and accountability.

### v3

@@ III-F System Design Considerations
III-G Slice Error Mitigation for Model Enhancement
@@ IV-A Hair Color Classifier Validation - Finding Edge Cases
$\bullet$ Does the model behave correctly on well-performing slices? One basic expected behavior for a hair color classification model is to catch hair features. At first glance of the Attribution Mosaic (Fig. 2  C), ML experts notice slice_22 and slice_5 on the left, which lie separately with others on the Attribution Mosaic.
@@
Their mosaic visualizations indicate gray hair patterns, which means the model uses the correct features, i.e., core features. By verifying the corresponding attribution heatmaps (Fig. 2  D), they confirm the correctness of this insight and annotate both slices as core feature with description Core: gray hair. Upon experts saving the annotation, AttributionScanner automatically propagates the annotation and provides a Spuriousness probability of each slice. With this guidance, the experts observe many slices lying on the right of Attribution Mosaic have higher Spuriousness probabilities (Fig. 2  C  E). Their mosaic visualizations do not show any hair patterns, leading to valid doubts that the model does not behave correctly on such slices. Besides, they also notice that the model has correct predictions on these slices (with $100\%$ prediction accuracy). This makes them worry that the model is largely biased by spurious features. Through investigation, they find the model mistakenly utilizes mouth and eyes to predict hair color for those top-performed slices, such as slice_25, slice_14, and slice_2 (Fig. 9).
$\bullet$ What underlying factors contribute to unexpected behaviors? By sorting the Slice Table by descending order of the propagated Spuriousness, ML experts notice more slices with high Spuriousness, such as slice_35, and are interested in understanding why such unexpected model behaviors happen. They switch Attribution Mosaic into the confusion matrix form (Fig. 2  E) and click on the name of “slice_35” to highlight this slice on the four sub-views and inspect explanations. They notice that images that lie in the “FN” group of this slice have wrong labels — they should be labeled as “not gray hair” rather than “gray hair”. They mark this issue as wrong labels and investigate its neighborhood slices on Attribution Mosaic, finding slice_2_FN also features wrong labels.
$\bullet$ Why slices underperform? After fully exploring those well-performed but biased slices, the experts wonder whether the underperformed slices are indeed issue-free or what contributes to their low performance. By sorting the Slice Table (Fig. 2  B) by ascending order of accuracy, they notice the model only achieves $72.41\%$ accuracy on slice_44 and find that the slice’s visualization only shows colorful patterns without any recognizable content. After checking the attribution heatmaps, they find that the model mistakenly uses image backgrounds to make hair color predictions (Fig. 9). This spurious correlation issue stands out, and they annotate this slice as “spurious” with the description “Spurious: backgrounds”. Similar issues also occur in its neighborhood slices via Attribution Mosaic, and AttributionScanner automatically propagates them with higher Spuriousness.
@@ IV-B Bird Category Classifier Validation - Detecting Bias
$\bullet$ Does the model exhibit bias? To answer this key question, experts start by investigating the underperformed slices. From the Slice Table (Fig. 7  A), they sort slices by ascending order of accuracy and select the worst-performed slice_38. The coordinated information provided by Attribution Mosaic and model attribution heatmaps highlights a spurious correlation problem—the model uses water backgrounds to classify birds (Fig. 10).
The experts annotate this slice as “spurious”, and AttributionScanner automatically propagates this annotation. They verify the propagation correctness on neighborhood slices and annotate slice_41 as “spurious” with “water backgrounds” (Fig. 10). Moreover, the experts identify an underperformed slice_32 that is not clustered with the current ones and is given a high Spuriousness possibility. They investigate it and verify that the model utilizes “Spurious feature: land backgrounds”, to predict bird classes. Through the Attribution Mosaic in the confusion matrix form (Fig. 7  D), they find such spurious correlations result in many false negatives (Fig. 7  D1), where the model uses land backgrounds to mistakenly predict many “waterbirds” as “landbirds”.
$\bullet$ Is the detected bias prevalent across all slices? Why or why not? ML experts are interested in determining whether the detected bias is pervasive throughout the dataset. By analyzing slices that are distant from the annotated ones in the Attribution Mosaic and are assigned with low Spuriousness possibilities, they discovered that the farthest neighbors, namely slice_34 and slice_5, correspond to core features. This suggests that the model can correctly capture bird regions in these slices (Fig. 7  B  C), which raises a follow-up “why” question. To understand in what circumstances when the model fails. They browse the original images from slice_32 (spurious feature) and slice_34 (core feature), respectively, as shown in Fig. 7  D and  C. They find slice_32 has very similar land backgrounds and very different birds, while on the other hand, the birds’ appearance of slice_34 is very consistent. This finding explains why this biased model can successfully capture the core features from slice_34 but fails at slice_32—greater similarities in the representation space indicate stronger features. In other words, core features in slice_34 are strong enough to support the model’s robustness against bias. Such insights are helpful in improving model robustness and have been further studied by ML experts [5].
@@
V Evaluation
V-A Quantitative Evaluation
$\bullet$ RCS. This metric quantifies the model’s reliance on core features while controlling for general noise robustness. It is defined as the ratio of the absolute gap between core and spurious accuracy, and the total possible gap for any model between core and spurious accuracy. Represented as $RCS=\frac{acc^{(C)}-acc^{(S)}}{2\times min(\alpha,1-\alpha)}$, where $\alpha=\frac{acc^{(C)}+acc^{(S)}}{2}$. RCS ranges from 00 to $1$, a higher value indicative of better model performance.
@@ V-C Experts Feedback
VI Conclusion
In this work, we present AttributionScanner, a novel human-in-the-loop model validation system empowered by metadata-free data slice finding and XAI techniques, which can support effective detection, explanation, and mitigation of potential errors and biases in vision models. AttributionScanner provides explainable data slices that reveal critical model problems, such as spurious correlations and mislabeled data. The effectiveness of this work is validated by the performance improvement and bias mitigation through both quantitative and qualitative evaluations. We hope this work can inspire future research in human-AI teaming to improve AI trustworthiness and accountability.
Acknowledgments
This research is supported in part by the National Science Foundation via grant No. IIS-2427770, and by the University of California (UC) Multicampus Research Programs and Initiatives (MRPI) grant and the UC Climate Action Initiative grant.
[…]
Appendix
This appendix is organized as follows:
  • Sec. A-I provides equations for calculating neighbor consistency of data features and model attributions, which are used to compare two feature representation spaces.
  • Sec. A-II provides equations for calculating slice attribution consistency.
  • Sec. A-III explains our interactive visual layer adjustment in the design of Attribution Mosaic.
  • Sec. A-IV details the Label Spreading Algorithm to complement the main paper.
  • Sec. A-V provides the discussion and future work.
A-I. Representation Space Comparison
In this section, we provide equations for calculating the local consistency of data features and model attributions, to complement the comparative study between the feature representation space (derived from the original feature vectors) and the attribution representation space (derived from the attribution-weighted feature vectors) as discussed in Sec. III-D(2) and Fig. 4 of the main paper.
Given a specific data sample $i$ and its neighbor sample $j$ on a representation space, we denote the original feature vector produced by model’s convolutional modules as $F_{*}$, and its GradCAM attribution mask as $M_{*}$. The data feature similarity between $i$ and $j$ is computed by:
$D_{sim}(F_{i},F_{j})=\frac{\langle F_{i},F_{j}\rangle}{\|F_{i}\|\cdot\|F_{j}\|},$ (A1)
which is the cosine similarity between $F_{i}$ and $F_{j}$, with a higher value indicative of higher similarity. The model attribution similarity between $i$ and $j$ is computed by:
$A_{sim}(M_{i},M_{j})=\frac{1}{W(M_{i},M_{j})},$ (A2)
where $W(M_{i},M_{j})$ represents the Wasserstein distance (also referred to as Earth Mover’s Distance) between the GradCAM attribution masks $M_{i}$ and $M_{j}$.
Then, for a specific sample $i$, the local consistency of data features is computed by the average $D_{sim}(F_{i},F_{*})$, and the local consistency of model attributions is computed by the average $A_{sim}(M_{i},M_{*})$, with higher values indicative of better local consistency, where * denotes the top 10 nearest neighbors of $i$ on the corresponding representation space.
A-II. Slice Attribution Consistency
To complement the data slice identification in Sec. III-D(2) of the main paper, we provide equations for calculating the attribution consistency of a data slice. Given a data slice $S_{i}$ containing $N$ attribution-weighted feature vectors, each denoted by $F_{W}$, the slice centroid is computed as:
$S_{i}^{c}=\frac{1}{N}\sum_{j=1}^{N}{F_{W}^{j}},F_{W}^{j}\in S_{i},$ (A3)
The attribution consistency of this slice is then calculated as:
$C_{S_{i}}=\frac{\sum_{j=1}^{N}{sim(F_{W}^{j},S_{i}^{c})}}{N},F_{W}^{j}\in S_{i},$ (A4)
where $sim(\cdot,\cdot)$ denotes cosine similarity.
A-III. Interactive Visual Layer Adjustment
In our Attribution Mosaic (refer to Sec. III-E of the main paper), we provide interactive visual layer adjustment to further support user exploration of our mosaic view. When hovering over a specific data slice, the corresponding slice mosaic would automatically expand and display on the top visual layer to avoid being covered by other mosaic panels. This automatic visual layer adjustment ensures clear visibility of all data slices in the mosaic view to facilitate user exploration.
A-IV. Label Spreading Algorithm
We detail the description of the Label Spreading Algorithm from [64] to complement Sec. III-E of the main paper.
The Label Spreading algorithm is a semi-supervised approach that propagates label information from a few annotated data points to unannotated ones by leveraging the similarity structure of the data. It assumes that similar data points should have similar labels and iteratively updates labels to enforce both local and global consistency. The key steps include:
1. Graph Construction. A graph is constructed where each data point is represented as a node, and the edge weights represent pairwise similarities between data points. The similarity between $x_{i}$ and $x_{j}$ can be defined based on different requirements, such as using a Gaussian kernel:
$w_{ij}=\exp\left(-\frac{\|x_{i}-x_{j}\|^{2}}{2\sigma^{2}}\right),$
where $\|x_{i}-x_{j}\|$ is the Euclidean distance, and $\sigma$ is the scaling parameter that controls the kernel width. In our case, we fit the graph to our attribution representation space and $x_{*}$ is the 2D coordinates of a sample on the space.
2. Graph Normalization. The weight matrix $W=[w_{ij}]$ is normalized to compute the transition matrix $T=D^{-1}W$, and $D$ is the diagonal degree matrix with $D_{ii}=\sum_{j}w_{ij}$.
3. Label Initialization. A label matrix $Y$ is initialized, where initial labels are binary values $\{0,1\}$ for annotated points and $-1$ for unannotated points.
4. Iterative Propagation. Labels are updated iteratively using:
$Y^{(t+1)}=\alpha TY^{(t)}+(1-\alpha)Y,$
where $Y^{(t)}$ is the label matrix at iteration $t$, $T$ is the transition matrix, $Y$ is the initial label matrix, and $\alpha\in[0,1]$ controls the balance between propagation and retention of initial labels.
5. Convergence. The algorithm iterates until $Y^{(t)}$ converges, meaning there is minimal change between successive iterations, ensuring the propagated labels are stable.
We use the off-the-shelf LableSpreading provided by scikit-learn [68] in our implementation with their default settings, ensuring simple deployment to apply to different cases.
A-V. Discussion and Future Work
@@
Propagation approaches for the hypothetical spuriousness.    We adopt the Label Propagation method [64] to automatically spread the slices’ spurious annotations from users to other slices as discussed in Sec. III-E (2). This is under the assumption that the distance between the slices corresponding to spurious correlations (e.g., water backgrounds) and the slices corresponding to core correlations (e.g., birds) is large. However, such assumptions can be invalid in other scenarios. For example, an image classification model trained on ImageNet [77] includes various different classes such as “husky dog” and “king penguin”. It is likely that two slices far away from each other are both corresponding to the core correlations. To address this, we plan to incorporate class similarities to the Label Propagation matrix to improve the precision of the generated hypothetical Spuriousness.